OpenStudy (anonymous):

Which equation will have the same solution as 3(2x – 1) + 4x = 13? A. 10x = 14 B. 9x = 14 C. 10x = 16 D. 9x = 16

OpenStudy (anonymous):

3(2x-1)+4x=13 6x-3+4x=13 10x-3=13 10x=16

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

No problem